OpenStudy (anonymous):

During a 4th of July weekend, 32 vehicles became trapped on the Sunshine Skyway Bridge while it was being repaved. A recent city ordinance decreed that only cars with 4 wheels and trucks with six wheels could be on the bridge at any given time. If there were 148 tires that needed to be replaced to due to damage, how many cars and trucks were involved in the incident? (2 points)

OpenStudy (anonymous):

Let there be x cars and y trucks involved in the accident. x+y =32 ...... Eq(1) 4x+6y = 148 => 2x+3y = 74 ..... Eq(2) 2* Eq(1) gives 2x+2y = 64 ...... Eq(3) Eq(2)- Eq.(3) gives 3y-2y = 74-64 => y =10 Use this value of y to solve for x in Equation 1, we get x = 22 So, there were 22 cars and 10 trucks involved.

OpenStudy (anonymous):

Alright, thanks!